Calculate the number of sodium ions, perchlorate ions, Cl atoms, and O atoms in 1.42 g of sodium perchlorate. Enter your answers in scientific notation.
1
Expert's answer
2019-10-02T08:01:47-0400

Find moles of NaClO4:


"n(NaClO_4)=\frac{m}{M}=\frac{1.42}{122.44}=0.0116 mol"

As one formula unit contains one sodium ion then n(Na+) =n(NaClO4) = 0.0116 mol



"N=N_A\times n"

"N(Na^+)=0.0116\times 6.022\times 10^{23}=6.99\times10^{21} ions"


As one formula unit contains one perchlorate ion then n(ClO4-) =n(NaClO4) = 0.0116 mol


"N(ClO_4^-)=0.0116\times 6.022\times10^{23} =6.99\times10^{21} ions"


As one formula unit contains one atom of chlorine then n(Cl) =n(NaClO4) = 0.0116 mol


"N(Cl) = 0.0116\times 6.022\times 10^{23} = 6.99\times10^{21} atoms"


As one formula unit contains 4 atoms of oxygen then n(O) =4*n(NaClO4) = 4*0.0116 mol=0.0464 mol


"N(O) = 0.0464\times 6.022\times10^{23} =2.79\times10^{22} atoms"